OpenStudy (anonymous):

need help please ... its my assignment ...derivatives of product and quotient ... pls answer this ... every 1 problem has 2 solution.product and a quotient ...Find y' if y=(x^3-2)/x2/3 ..answer it please

myininaya (myininaya):

\[y=\frac{x^3-2}{x^{\frac{2}{3}}}\] ? if this is right, then it is silly to use the quotient rule/product rule

myininaya (myininaya):

\[y=x^{3-\frac{2}{3}}-2x^{\frac{-2}{3}}\] \[y=x^{\frac{7}{3}}-2x^\frac{-2}{3}\] just use power rule \[y'=\frac{7}{3}x^{\frac{7}{3}-1}-2*\frac{-2}{3}x^{\frac{-2}{3}-1}\]

myininaya (myininaya):

\[y'=\frac{7}{3}x^\frac{4}{3}+\frac{4}{3}x^\frac{-5}{3}\]